SOLUTION: find the equation of the line that is perpendicular to the line y=2x-3 and passes through the point (-2,1)
Answer by Alan3354(69443)

SOLUTION: find the equation of the line that is perpendicular to the line y=2x-3 and passes through the point (-2,1)
------------
The slope of y=2x-3 is 2. (Did you know that?)
The slope m of lines perpendicular is the negative inverse, = -1/2
------
y-y1 = m*(x-x1)
y-1 = (-1/2)*(x+2)